# Let the Void Be Void: Robust Open-Set Semi-Supervised Learning via Selective Non-Alignment

### v3

@@ B. Prototype Generation
$w_{lk}=\gamma\cdot n_{lk},\quad w_{uk}=r_{u}\cdot n_{uk}$ (10)
@@
$w_{lk}^{\text{norm}}=\frac{w_{lk}}{w_{lk}+w_{uk}},\quad w_{uk}^{\text{norm}}=\frac{w_{uk}}{w_{lk}+w_{uk}}$ (11)
@@
$\mu_{k}=w_{lk}^{\text{norm}}\cdot\mu_{lk}+w_{uk}^{\text{norm}}\cdot\mu_{uk}$ (12)
@@ E.2 Training Details
For CIFAR-100, we used a 1-layer MLP with ReLU for the closed-set classifier $\mathrm{CC}(\cdot)$ and a 2-layer MLP with ReLU for the outlier detector $\mathrm{OD}(\cdot)$, both with a hidden size of 1024. The projection layer in the SNA module consists of a 1-layer MLP with ReLU, with a hidden size of 256 and a projection size of 128. For CIFAR-10, $\mathrm{CC}(\cdot)$ uses a smaller hidden size of 128, and the projection layer’s hidden size is also reduced to 128. For ImageNet30 and TinyImageNet, $\mathrm{CC}(\cdot)$ uses a hidden size of 256, and the projection layer uses a hidden size of 256, while the projection size remains 128. For CIFAR-100, the model was trained with a batch size of 256, $\gamma=2$, and 1024 iterations per epoch for 256 epochs using 4 GPUs (RTX 3090). For CIFAR-10, training was conducted on a single GPU (RTX 3090) with a batch size of 64, $\gamma=4$, over 256 epochs. For ImageNet30, we used 4 GPUs (RTX 3090) with a batch size of 64, $\gamma=2$, for 512 epochs. For TinyImageNet, the model was trained on a single GPU (H100) with a batch size of 32 and $\gamma=6$ for 512 epochs. All experiments used the SGD optimizer with momentum 0.9, and a cosine annealing learning rate schedule starting from 0.03. Weight decay was set to 0.001 for CIFAR-100 and 0.0005 for all other datasets. All experiments were run with three random seeds: 0, 1, and 2. Additional hyperparameter details are available in the experimental configuration files provided in the code repository.
@@
1:Input: Labeled dataset $\mathcal{D}_{l}$, labels $y\in\{1,\dots,K\}$, unlabeled dataset $\mathcal{D}_{u}$, data augmentations $\mathcal{T}_{t}$ for $t\in\{w,w^{\prime},s\}$, feature extractor(Shared backbone) $\mathcal{F}(\cdot)$, classifiers $g_{CC}$ and OOD detector $g_{OD}$, output heads $CC(\cdot)$, $OD(\cdot)$, projection head $proj(\cdot)$, class prototypes $\mu$, iteration count ItIt
2:
3:for $i=1$ to ItIt do
4:  /* Training Data Preparation */
5:  Sample labeled batch $X=\{(x_{i},y_{i})\}_{i=1}^{B}$
6:  Sample unlabeled batch $\mathcal{U}=\{u_{i}\}_{i=1}^{\gamma B}$
7:  Apply augmentations:
8:    $x_{i}^{t}=\{\mathcal{T}_{w}(x_{i}),\mathcal{T}_{w^{\prime}}(x_{i}),\mathcal{T}_{s}(x_{i})\}$
9:    $u_{i}^{t}=\{\mathcal{T}_{w}(u_{i}),\mathcal{T}_{w^{\prime}}(u_{i}),\mathcal{T}_{s}(u_{i})\}$
10:  /* Forward Pass */
11:  $h_{i}^{t}=\mathcal{F}(x_{i}^{t})$,  $h_{u,i}^{t}=\mathcal{F}(u_{i}^{t})$
12:  $z_{i}^{t}=proj(h_{i}^{t})$,  $z_{u,i}^{t}=proj(h_{u,i}^{t})$
13:  $p_{i}^{t}=CC(h_{i}^{t})\in\mathbb{R}^{K}$,  $p_{u,i}^{t}=CC(h_{u,i}^{t})$
14:  $\varphi_{i}^{t}=OD(h_{i}^{t})\in\mathbb{R}^{2K}$ $\triangleright$ $\varphi_{i,k}^{t}=\{\varphi_{i\text{-}ID,k}^{t},\ \varphi_{i\text{-}OOD,k}^{t}\}$
15:  $\varphi_{u,i}^{t}=OD(h_{u,i}^{t})\in\mathbb{R}^{2K}$
16:  /* Compute Losses */
17:  $\mathcal{L}_{USNA}=\text{Unlabeled\_SNA\_Loss}(z_{u,i}^{w})$ $\triangleright$ Eq. 2
18:  $\mathcal{L}_{IA}=\text{Instance-wise\_Alignment\_Loss}(z_{i}^{w})$ $\triangleright$ Eq. 4
19:  $\mathcal{L}_{PA}=\text{Prototype-based\_Alignment\_Loss}(z_{i}^{w})$ $\triangleright$ Eq. 5
20:  $\mathcal{L}_{SNA}=\lambda_{USNA}\mathcal{L}_{USNA}(\mathcal{U})+\lambda_{IA}\mathcal{L}_{IA}(X)+\lambda_{PA}\mathcal{L}_{PA}(X)$
21:  $\mathcal{L}_{CC}=\mathcal{L}_{x}(X)+\lambda_{u}\mathcal{L}_{u}(\mathcal{U})$
22:  $\mathcal{L}_{OD}=\mathcal{L}_{od}^{OVA}(X)+\lambda_{em}\mathcal{L}_{od}^{em}(\mathcal{U})+\lambda_{SOCR}\mathcal{L}_{od}^{SOCR}(\mathcal{U})+\lambda_{neg}\mathcal{L}_{od}^{neg}(\mathcal{U})$
23:  $\mathcal{L}_{Total}=\lambda_{CC}\mathcal{L}_{CC}+\lambda_{OD}\mathcal{L}_{OD}+\lambda_{SNA}\mathcal{L}_{SNA}$
24:  Update model with loss $\mathcal{L}$
25:  if current iteration is a prototype update step then
26:   Call Prototype_generation($z_{i}^{w},z_{u,i}^{w}$)
27:  end if
28:end for
29:
30:Prototype_generation($z_{i}^{w}$, $z_{u,i}^{w}$):
31:for $k=1$ to $K$ do
32:  $\mu_{lk}=\text{mean}(z_{i}^{w}\mid y_{i}=k)$
33:  $\mu_{uk}=\text{mean}(z_{u,i}^{w}\mid$ reliable ID for class $k)$
34:  $\mu_{k}=w_{lk}^{norm}\cdot\mu_{lk}+w_{uk}^{norm}\cdot\mu_{uk}$ $\triangleright$ Eq. 12
35:end for

### v4

@@ B. Prototype Generation
$w_{lk}=\gamma\cdot n_{lk},\quad w_{uk}=r_{u}\cdot n_{uk}$
@@
$w_{lk}^{\text{norm}}=\frac{w_{lk}}{w_{lk}+w_{uk}},\quad w_{uk}^{\text{norm}}=\frac{w_{uk}}{w_{lk}+w_{uk}}$
@@
$\mu_{k}=w_{lk}^{\text{norm}}\cdot\mu_{lk}+w_{uk}^{\text{norm}}\cdot\mu_{uk}$
@@ E.2 Training Details
For CIFAR-100, we used a 1-layer MLP with ReLU for the closed-set classifier $\mathrm{CC}(\cdot)$ and a 2-layer MLP with ReLU for the outlier detector $\mathrm{OD}(\cdot)$, both with a hidden size of 1024. The projection layer in the SNA module consists of a 1-layer MLP with ReLU, with a hidden size of 256 and a projection size of 128. For CIFAR-10, $\mathrm{CC}(\cdot)$ uses a smaller hidden size of 128, and the projection layer’s hidden size is also reduced to 128. For ImageNet30 and TinyImageNet, $\mathrm{CC}(\cdot)$ uses a hidden size of 256, and the projection layer uses a hidden size of 256, while the projection size remains 128. For CIFAR-100, the model was trained with a batch size of 256, $\gamma=2$, and 1024 iterations per epoch for 256 epochs using 4 GPUs (RTX 3090). For CIFAR-10, training was conducted on a single GPU (RTX 3090) with a batch size of 64, $\gamma=4$, over 256 epochs. For ImageNet30, we used 4 GPUs (RTX 3090) with a batch size of 64, $\gamma=2$, for 512 epochs. For TinyImageNet, the model was trained on a single GPU (H100) with a batch size of 32 and $\gamma=6$ for 512 epochs. All experiments used the SGD optimizer with momentum 0.9, and a cosine annealing learning rate schedule starting from 0.03. Weight decay was set to 0.001 for CIFAR-100 and 0.0005 for all other datasets. All experiments were run with three random seeds: 0, 1, and 2. Additional hyperparameter details are available in the experimental configuration files provided in the code repository. Code is available at https://github.com/yourimchoi/skipalign
@@
1:Input: Labeled dataset $\mathcal{D}_{l}$, labels $y\in\{1,\dots,K\}$, unlabeled dataset $\mathcal{D}_{u}$,
2:   data augmentations $\mathcal{T}_{t}$ for $t\in\{w,w^{\prime},s\}$,
3:   feature extractor(Shared backbone) $\mathcal{F}(\cdot)$, classifiers $g_{CC}$ and OOD detector $g_{OD}$,
4:   output heads $CC(\cdot)$, $OD(\cdot)$, projection head $proj(\cdot)$,
5:   class prototypes $\mu$, iteration count ItIt
6:for $i=1$ to ItIt do
7:  /* Training Data Preparation */
8:  Sample labeled batch $X=\{(x_{i},y_{i})\}_{i=1}^{B}$
9:  Sample unlabeled batch $\mathcal{U}=\{u_{i}\}_{i=1}^{\gamma B}$
10:  Apply augmentations:
11:    $x_{i}^{t}=\{\mathcal{T}_{w}(x_{i}),\mathcal{T}_{w^{\prime}}(x_{i}),\mathcal{T}_{s}(x_{i})\}$
12:    $u_{i}^{t}=\{\mathcal{T}_{w}(u_{i}),\mathcal{T}_{w^{\prime}}(u_{i}),\mathcal{T}_{s}(u_{i})\}$
13:
14:  /* Forward Pass */
15:  $h_{i}^{t}=\mathcal{F}(x_{i}^{t})$,  $h_{u,i}^{t}=\mathcal{F}(u_{i}^{t})$
16:  $z_{i}^{t}=proj(h_{i}^{t})$,  $z_{u,i}^{t}=proj(h_{u,i}^{t})$
17:  $p_{i}^{t}=CC(h_{i}^{t})\in\mathbb{R}^{K}$,  $p_{u,i}^{t}=CC(h_{u,i}^{t})$
18:  $\varphi_{i}^{t}=OD(h_{i}^{t})\in\mathbb{R}^{2K}\quad\triangleright\ \varphi_{i,k}^{t}=\{\varphi_{i\text{-}ID,k}^{t},\ \varphi_{i\text{-}OOD,k}^{t}\}$
19:  $\varphi_{u,i}^{t}=OD(h_{u,i}^{t})\in\mathbb{R}^{2K}$
20:
21:  /* Compute Losses */
[…]
25:  $\mathcal{L}_{SNA}=\lambda_{USNA}\mathcal{L}_{USNA}(\mathcal{U})+\lambda_{IA}\mathcal{L}_{IA}(X)+\lambda_{PA}\mathcal{L}_{PA}(X)$
26:  $\mathcal{L}_{CC}=\mathcal{L}_{x}(X)+\lambda_{u}\mathcal{L}_{u}(\mathcal{U})$
27:  $\mathcal{L}_{OD}=\mathcal{L}_{od}^{OVA}(X)+\lambda_{em}\mathcal{L}_{od}^{em}(\mathcal{U})+\lambda_{SOCR}\mathcal{L}_{od}^{SOCR}(\mathcal{U})$ + $\lambda_{neg}\mathcal{L}_{od}^{neg}(\mathcal{U})$
28:   $\mathcal{L}_{od}^{OVA}(\mathcal{X})=\sum_{i=1}^{B}\sum_{k=1}^{K}-\mathds{1}(y_{i}=k)\log\sigma(s_{k}^{\text{ID}})-\mathds{1}(y_{i}\neq k)\log\sigma(s_{k}^{\text{OOD}})$
29:   $\mathcal{L}_{od}^{em}(\mathcal{U})=-\frac{1}{\gamma B}\sum_{i=1}^{\gamma B}\sum_{k=1}^{K}\{\varphi_{i,k}^{ID}\log(\varphi_{i,k}^{ID})+\varphi_{i,k}^{OOD}\log(\varphi_{i,k}^{OOD})\}$
30:   $\mathcal{L}_{od}^{SOCR}(\mathcal{U})=\frac{1}{\gamma B}\sum_{i=1}^{\gamma B}\sum_{k=1}^{K}|s_{i,k,\mathcal{T}_{w}}^{ID}-s_{i,k,\mathcal{T}_{w^{\prime}}}^{ID}|^{2}$
31:   $\mathcal{L}_{od}^{neg}(\mathcal{U})=-\frac{1}{\gamma B}\sum_{i=1}^{\gamma B}\frac{1}{\sum_{k=1}^{K}\mathds{1}(\varphi_{i,k}^{ID}<\eta_{neg})}\sum_{k=1}^{K}\mathds{1}(\varphi_{i,k}^{ID}<\eta_{neg})\log(1-\varphi_{i,k}^{ID})$
32:  $\mathcal{L}_{Total}=\lambda_{CC}\mathcal{L}_{CC}+\lambda_{OD}\mathcal{L}_{OD}+\lambda_{SNA}\mathcal{L}_{SNA}$
33:  Update model with loss $\mathcal{L}$
34:  if current iteration is a prototype update step then
35:   Call Prototype_generation($z_{i}^{w},z_{u,i}^{w}$)
36:  end if
37:end for
38:
39:Prototype_generation($z_{i}^{w}$, $z_{u,i}^{w}$):
40:for $k=1$ to $K$ do
41:  $\mu_{lk}=\text{mean}(z_{i}^{w}\mid y_{i}=k)$
42:  $\mu_{uk}=\text{mean}(z_{u,i}^{w}\mid\text{reliable ID for class }k)$
[…]
44:end for